OpenStudy (twistrose):

What is the value of x? 6(−3−x)−2x=14

OpenStudy (leenathan):

twistey

OpenStudy (leenathan):

ok start to use pemdas

OpenStudy (twistrose):

....ok how am i gonna use that when x is in parentisis and i have to find the value of x

OpenStudy (peachpi):

is that put in right? there are two equal signs...

OpenStudy (twistrose):

no..it doubled again

OpenStudy (peachpi):

what's it supposed to be?

OpenStudy (peachpi):

oh nvm

OpenStudy (peachpi):

start by distributing the 6

OpenStudy (peachpi):

6(-3-x) = ?

OpenStudy (twistrose):

ok so: -18-6x

OpenStudy (twistrose):

@peachpi

OpenStudy (peachpi):

yes. so now the equation is\[-18-6x-2x=14\]You can combine the -6x and -2x to get\[-18-8x=14\]Can you solve from here?

OpenStudy (twistrose):

Ii think so